OpenStudy (anonymous):

Find the equation for the line through (2,6) and perpendicular to y = -5/4x+1

OpenStudy (anonymous):

Helllpppp

OpenStudy (anonymous):

Wait for a moment. I ´m solving that. but i will give u a tip. The angular coefficient must be m1=-1/m2

OpenStudy (anonymous):

ok thanks! Let me know what you get

OpenStudy (anonymous):

So, we should have m1=-5/4/m2

OpenStudy (anonymous):

yes. Now what

OpenStudy (anonymous):

Use the formalae of general line, the equation y=ax+b. and substitute the x and y for the values taken for the pair (2,6).

OpenStudy (anonymous):

yes, i got -5/4x+17/2 is that right?!

OpenStudy (anonymous):

????

OpenStudy (anonymous):

So, substituting in the equation you got 6=-8/5+b; -b=-8/5-6

OpenStudy (anonymous):

And then we got b = 38/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Using newly the equation y=ax+b we have y=-4x/5-38/5

OpenStudy (anonymous):

the answer similar to the in multiple choice is -4/5x+38/5....

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Don´t i deserve a medal?